This is a problem to determine the width and length of a rectangle where the length is 45 feet more than the width. The area is 9000 feet. What formula is used to determined the width and length. We tried and tried to figure out, but can't do it. Please help
Answer by mananth(16946)

width = x
length = x+45

Area = L * W
Area = x(x+45)
x(x+45)=9000
x^2+45x=9000
x^2+45x-9000=0
x^2+120x-75x-9000=0
x(x+120)-75(x+120)=0
(x+120)(x-75)=0
x= -120, x=75
width = 75 feet
length = 75+45=120 feet